Part H Constants What is the charge per unit length on the outer surface of the tube? Express your answer in terms of the given quantities and appropriate constants. A very long conducting tube (hollow cylinder) has inner radius a and outer radius b. It carries charge per unit length-a where α is a positive constant with units of C/m A line of charge lies along the axis of the tube. The line of charge has charge per unit length +a Previous Answers Request Answer Submit incorrect; Try Again; 4 attempts remaining
Answer #1

Charge per unit length of long conducting tube ( hollow cylinder) = - \alpha C/m

Chage per unit length of line charger = \alpha C/m

As we know that charge placed inside a hollow conducting shell gets uniformly distributed at outer surface of tube. Although physically line charge per unit length is present inside a hollow tube but it appears ( it's effect) to be distributed on the surface.

So net charge on outer per unit length of outer surface of tube

= -\alpha+ \alpha = 0

Net charge on outer surface = 0